# The Future of Large Language Model Pre-training is Federated

### v1

@@ Abstract
Generative pre-trained large language models (LLMs) have demonstrated impressive performance over a wide range of tasks, thanks to the unprecedented amount of data they have been trained on. As established scaling laws indicate, LLMs’ future performance improvement depends on the amount of computing and data sources we can leverage for pre-training. Federated learning (FL) has the potential to unleash the majority of the planet’s data and computational resources, which are underutilized by the data-center-focused training methodology of current LLM practice. Our work presents a robust, flexible, reproducible FL approach that enables large-scale collaboration across institutions to train LLMs. This would mobilize more computational and data resources while matching or potentially exceeding centralized performance. We further show the effectiveness of the federated training scales with model size and present our approach for training a billion-scale federated LLM using limited resources. This will help data-rich actors to become the protagonists of LLMs pre-training instead of leaving the stage to compute-rich actors alone.
@@ 1 Introduction
The impressive performance of generative pre-trained large language models (LLMs) and their multi-modal derivations largely owes to their capacity to learn representations at scale [1]. Thus, a very small number of well-resourced tech companies and institutions are using increasingly powerful computing facilities in the race to scale up LLMs and dataset sizes to achieve state-of-the-art (SOTA) performance. The thousands of hours of training to convergence on thousands of specialized and well-connected hardware accelerators in a single data center incur a high energy and monetary cost [2]. Distributing training across multiple data centers in sparse geographical locations, for those companies who could afford such a thing, would drive the cost even higher due to communication overheads [3, 4].
@@
Hoffmann et al. [5] showed that the effective performance improvement of increasingly large LLMs requires increasingly extensive training datasets. Most of the world’s data sources are unevenly distributed among private actors who often don’t want to share it, even if the data regulations of their respective jurisdictions permit such sharing. Since no organization independently owns the rights to a sufficient amount of text data, the multi-terabyte datasets used in these procedures must be obtained from publicly available sources. These may include materials potentially protected by intellectual property laws [6] or be otherwise problematic [7, 8, 9, 10]. The potential means of collecting non-public data may require cutting independent deals with data providers [11] or taking model training directly to these private data sources. The further improvement of LLMs will require more extensive high-quality language data than what has been estimated [12] to be publicly available in the future, as early as $2026$. Given the limits on dataset growth coupled with near-unbounded model growth, increases in memorization and data leakage amongst current LLMs [13, 14] have started to become evident.
@@
The next generation of LLMs and foundation models (FMs) will benefit from effectively leveraging more data and computational sources than the centralized paradigm currently makes available. Therefore, we argue that making FMs future even brighter requires shifting the dominant training paradigm to a collaborative federated approach, where nodes control considerable but not immense computational or data resources according to their abilities. Using federated learning (FL), we can expand the quality of our models by gaining access to previously untapped data and computational sources [15]. This will, in turn, allow us to increase the size of models that can be efficiently trained compared to the centralized paradigm and to avoid both memorization and data leakage.
@@
As shown in Stich [16], Lin et al. [17], McMahan et al. [18], FL can relax the synchronization requirements of stochastic gradient descent (SGD) to accommodate such poorly connected nodes. Also, Douillard et al. [19], Liu et al. [20], Douillard et al. [21] showed that Local SGD could substantially reduce the communication overhead of training LLMs in data center settings with homogeneous and heterogeneous computational nodes. Advancing beyond their useful independent insights, we argue for an entirely federated approach to LLM training that can automatically balance the workload of multiple actors and privately merge the knowledge derived from their local datasets without divulging them directly to any other participant.
@@
We are the first [22] to succeed in generative pre-training a billion-scale model in a heterogeneous federated setting. This work presents a complete system for pre-training federated LLMs in a collaborative, reproducible, and scalable fashion that can address the main challenges for a flexible execution in arbitrary FL cross-silo settings. It is built on the open-source FL framework Flower [23] and shall be released to the public after light refactoring. Furthermore, we make this available as a training recipe with a fully transparent experimental configuration accompanying our results. We emphasize that federated generative pre-training can be done with affordable hardware configurations by interlinking single nodes containing $1$-$8$ GPUs from standard cloud providers rather than renting entire data centers at once. We demonstrate that pre-training LLMs can be democratized through FL technologies, including data-rich actors independent of their connectivity and computing resources.
@@
Between the numerous insights discovered in developing this novel federated system, we show, for the first time, that larger federated LLMs find a consensus across clients more easily than their smaller versions, contrary to our expectations. Our results show that:
@@
    Under our system, larger models require less frequent communication while receiving a greater boost in generalization performance than smaller ones.
@@
    By utilizing our execution engine, Pollen [24], to balance loads across computational nodes, we can establish a training pipeline that is faster and more robust than its centralized counterpart.
@@ 2 The Landscape of LLM Training
Generative pre-trained large language models (LLMs) have demonstrated powerful performance across various natural language processing tasks, leading to rapid and widespread adoption. They are trained on massive corpora of text, incorporating mixtures of low-quality web-scraped data and high-quality curated datasets [25, 26]. Recently, several institutions have released their pre-trained LLMs, either in the closed form, such as GPT-4 [26], Chinchilla [5], and Gemini [27], or the open-sourced form, such as BLOOM [28], LLaMa [29, 30], and Falcon [31]. The scaling laws identified by Kaplan et al. [1], Hoffmann et al. [5] dictate that model size and dataset size should be increased in equal measure to improve model performance best. These suggest a future race between entities interested in developing state-of-the-art LLMs to grab as many compute and data sources as possible. The promising direction in which this field is headed can become even more luminous by gaining the trust of private entities, which possess an unprecedented breadth of knowledge and computing resources [15].
In Section 2.1, we will describe the current landscape for generative pre-training of LLMs with particular attention to the techniques for centralized distributed training. We will discuss federated learning (FL) and its impact as a communication efficient technique in Section 2.2. Section 2.3, for completeness, presents the effort the community has put into matching FL with LLMs fine-tuning. However, we highlight that our work tackles the far more challenging problem of federated pre-training of LLMs.
@@ 2.1 Centralized Distributed Optimization
LLM pre-training is based on two denoising pillars: leveraging huge batch sizes and very long contexts, i.e., sequence length of a single input sample. The combination of model and batch sizes forces training LLMs to scale SGD beyond the confines of a single GPU. Thus, it is often necessary to process more training samples in parallel and to split the model across GPUs if it is incapable of fitting within the memory of one GPU. In the following, we present a subset of optimizations and techniques that have been proposed to make this challenging training recipe possible. As we will see in Section 2.2, Federated Learning follows as a natural next step in the sequence of distributed training optimizations previously adopted by the field. Given the importance of such optimizations even for local training in federated learning, our work supports most of them.
@@ 2.1.1 Data and Model Parallelism
The number of trainable parameters and the size of the datasets make LLM training very sensitive to the stochastic fluctuations of the optimizer used, thus requiring a strong and robust regularization achieved by the denoising properties of very large batch sizes. To enable training with larger effective batch size, Data Parallelism replicates the model $N_{d}$ times across different devices and feeds each replica with training batches in parallel; thus, each replica produces the gradients of its training micro-batch locally and accumulates them via simple summation. This accumulation is eventually followed by a synchronization step where these gradients are averaged and applied to the model. Since the batch size of an ML algorithm plays a crucial role in the convergence of the training procedure [32, 33], this synchronization step happens after each replica has accumulated sufficient micro-batches to match the desired true batch size. Modern data parallelism implementations such as the one used by PyTorch Distributed [34] use the Ring AllReduce algorithm popularized by Horovod [35], implemented with low-level collective communication libraries like NCCL, to reduce the gradients across replicas. Ring AllReduce scales proportionally to the amount of data being transmitted and the latency of the slowest connection between workers in the ring, making it highly sensitive to the network topology connecting GPUs.
In the case of LLM training, the model size often mandates computing gradients over batch sizes larger than what GPUs can support, even with Data Parallelism. Thus, they are forced to use a micro-batch size dictated by the number of training samples that fit in the GPU, which is, in turn, limited by the model’s VRAM occupancy and activations. These factors result in a complex interplay between the model’s memory consumption and the training pipeline’s efficiency. This is exacerbated when considering how models may be partitioned across GPUs to reduce their VRAM consumption.
For sufficiently large models, the parameters must be split across GPU workers so that they fit in VRAM. Traditionally, this was achieved via Model Parallelism (MP) [36, 37], splitting individual tensors and their computation vertically across GPUs and communicating as needed. While model parallelism can reduce the per-GPU memory requirements linearly by the model parallelism degree $N_{m}$, it induces communication overheads that are tolerable in a single-machine context with high inter-GPU bandwidth but cannot scale to multi-machine contexts.
@@ 2.1.2 Fully Sharded Data Parallelism
An alternative approach is to shard the model into equally-sized units amongst GPUs, with units potentially containing multiple layers, and then materialize the units, as necessary, to compute the activations during the forward pass via collective communication. This form of fully-sharded data parallelism [38, 39] reduces memory consumption linearly in the data-parallelism degree $N_{d}$ while increasing communication by $1.5$$\times$ compared to standard data parallelism. It is also possible to combine this methodology with other techniques for reducing memory consumption, such as model parallelism, activation checkpointing [40], or CPU offloading [41]. Activation checkpointing functions by recomputing activations during the backward pass rather than saving them in memory. CPU offloading refers to offloading either data to system RAM or operations to the CPU.
Given the memory requirements, it is important to consider the minimum number of GPUs necessary to train one model with reasonable efficiency, e.g., without extreme CPU offloading. This provides a lower bound on the hardware that an actor requires to participate in any distributed training of an LLM and is generally determined by the model size and the micro-batch size. By accounting for each actor’s independent resources and manipulating the amount of local computation they do together with the micro-batch size, our method can relax the lower bound within reasonable limits, thus allowing even actors with weak hardware to participate in distributed training.
@@ 2.1.3 Bottlenecks for generative pre-training of LLMs
Similarly, hardware accelerators with enough memory and throughput to support LLM training are scarce and increasingly unaffordable to all actors except for the best-funded organizations. Hundreds to thousands of such accelerators are required with extremely high monetary costs for training and inference [28, 43]. Moreover, such accelerators must be largely uniform in specifications to avoid stragglers and need to be extremely well-connected both within a computational node and across nodes due to the synchronization requirements of data parallel or fully shared data parallel SGD [34, 38, 39]. These constraints pose significant system challenges that can be mitigated by extremely expensive configurations or by developing heterogeneity-aware pipelines that exploit the resources available despite their heterogeneity. The difficulties described above scale with model size, as splitting the model across the memory of several GPUs further increases communication demands [38, 39] and the optimization gaps.
@@ 2.1.4 Mitigation of LLMs demands
The wider AI community has gone to great lengths to circumvent the high requirements of LLM utilization and thus expand the data pool we can draw upon. Such endeavors have widely focused on locally running efficient inference with pre-trained model weights [29, 30], quantization [44], or parameter-efficient fine-tuning [45]. The recent work of Borzunov et al. [46] proposes Petal, which enables wide-scale collaboration for inference and parameter-efficient fine-tuning over the Internet by joining the resources of multiple parties. Their work assumes that clients provide inference jobs or data for fine-tuning, and servers execute the LLM inference/fine-tuning in a distributed pipeline parallel fashion. To enable efficient fine-tuning, their system makes clients responsible for holding the trainable parameters, with servers merely running forward passes and returning gradients for the pre-trained weights they store.
We argue that while methods exploiting pre-trained weights are highly beneficial to the wider community, they do not resolve the bottleneck of pre-training. Thus, the community is bound to the decisions made by actors capable of training LLMs regarding their structure and the data they use, and they may suffer downstream consequences for such reliance. This may happen since the pre-trained model, further fine-tuned for a specific downstream task, is the dominant upper bound for the downstream model’s performances. Thus, we propose developing systems capable of the distributed pre-training of LLMs that can accommodate the variety of hardware and data available in the AI community. This may include training from scratch if the model’s size permits, starting from pre-initialized weights and retraining the entire model instead of only a limited subset of parameters.
@@ 2.2 Federated Learning and Local SGD
Traditional machine learning involves using a central server that hosts the machine learning models and all the data in one place. In contrast, in Federated Learning (FL) [18] frameworks, client devices collaboratively learn a shared global model using their local compute and data.
FL aims to collaboratively learn a global model while keeping private data on the device. FL training occurs over multiple communication rounds. During each round, a fraction of the clients are selected and receive the global model from the server. Those selected clients then perform local training with their local data before sending the updated models back to the central server. Finally, the central server aggregates these updates via averaging [18] into a pseudo-gradient. It then uses a federated optimizer [47] to update its model based on the pseudo-gradient, creating a new global model. Then, this three-stage process is repeated.
@@
Despite these advantages, FL comes with two major challenges in the form of data and systems heterogeneity [50]. Data heterogeneity refers to the tendency of naturally generated and partitioned data to fail the IID assumption, which is common in centralized ML optimization. Systems heterogeneity refers to the ability of client hardware to vary in terms of computational ability, communication efficiency, or availability [51]. Both forms of heterogeneity are highly relevant for the distributed training of LLMs. Data heterogeneity may arise from participants holding texts in different languages, belonging to different genres, or varying in complexity. Systems heterogeneity is primarily present in the computational ability of the GPUs held by a specific client, their VRAM, and number, as well as the communication efficiency of said client.

### v2

@@ Abstract
Generative pre-trained large language models (LLMs) have demonstrated impressive performance over a wide range of tasks, thanks to the unprecedented amount of data they have been trained on. As established scaling laws indicate, LLMs’ future performance improvement depends on the amount of computing and data sources they can leverage for pre-training. Federated learning (FL) has the potential to unleash the majority of the planet’s data and computational resources, which are underutilized by the data-center-focused training methodology of current LLM practice. Our work presents a robust, flexible, reproducible FL approach that enables large-scale collaboration across institutions to train LLMs. We propose a scalable deployment system called Photon to enable the investigation and development of this new training paradigm for LLM pre-training. We show that Photon can be used by organizations interested in collaborating with their private data sources and computational resources for pre-training LLMs with billions of parameters. This paradigm would mobilize more computational and data resources while matching or potentially exceeding centralized performance. We further show the effectiveness of the federated training scales with model size and present our approach for training a billion-scale federated LLM using limited resources. Finally, we show that LLM training is highly resilient to the classical challenges of federated statistical and hardware heterogeneity. Furthermore, we show that convergence is robust to partial participation, opening the avenue for compute-efficient collaborative training. Photon will help data-rich actors to become the protagonists of LLMs pre-training instead of leaving the stage to compute-rich actors alone.
@@ 1 Introduction
The impressive performance of generative pre-trained large language models (LLMs) and their multi-modal derivations largely owes to their capacity to learn representations at scale [2]. Thus, a handful of well-resourced tech companies and institutions are using increasingly powerful computing facilities in the race to scale up LLMs and dataset sizes to achieve state-of-the-art (SOTA) performance. The thousands of hours of training to convergence on thousands of specialized and well-connected hardware accelerators in a single data center incur a high energy and monetary cost [3]. Distributing training across multiple data centers in sparse geographical locations, for those companies who could afford it, would drive the cost even higher due to communication overheads [4, 5].
@@
Hoffmann et al. [6] showed that the effective performance improvement of increasingly large LLMs requires increasingly extensive training datasets. Most of the world’s data sources are unevenly distributed among private actors who often don’t want to share it, even if the data regulations of their respective jurisdictions permit such sharing. Since no organization independently owns the rights to a sufficient amount of text data, the multi-terabyte datasets used in these procedures must be obtained from publicly available sources. These may include materials potentially protected by intellectual property laws [7] or be otherwise problematic [8, 9, 10, 11]. The potential means of collecting non-public data may require arranging independent deals with data providers [12] or taking model training directly to these private data sources. It is anticipated that, by $2026$, the amount of high-quality data that LLMs will require to continue to improve their performance will exceed what is publicly available [13]. Where datasets’ sizes become constrained, but model growth is unbounded, this can lead to memorization and data leakage amongst current LLMs [14, 15].
@@
The next generation of LLMs and foundation models (FMs) will benefit from effectively leveraging more data and computational resources than the centralized paradigm currently makes available. Therefore, we argue that making FMs’ future even brighter requires shifting the dominant training paradigm to a collaborative federated approach, where nodes control considerable (but not immense) computational or data sources according to their abilities. Using federated learning (FL), we can expand the quality of our models by gaining access to previously untapped data and computational sources [16]. This broader data availability will, in turn, allow us to increase the size of models that can be efficiently trained compared to the centralized paradigm and to avoid both memorization and data leakage.
@@
As shown in previous works [17, 18, 19], FL can relax the synchronization requirements of stochastic gradient descent (SGD) to accommodate such poorly connected nodes. Also, more recent works [20, 21, 22] showed that Local SGD could substantially reduce the communication overhead of training LLMs in data center settings with homogeneous and heterogeneous computational nodes. Advancing beyond their valuable independent insights, we argue for an entirely federated approach to LLM training that can automatically balance the workload of multiple actors and privately merge the knowledge derived from their local datasets without divulging them directly to other participants.
@@
We are the first [1] to fully utilize a practical and production-ready geographically distributed federation of heterogeneous devices and data sources for reproducible generative pre-training of a billion-scale LLM. This work presents Photon, a complete system for pre-training federated LLMs in a collaborative, reproducible, and scalable fashion that can address the main challenges for a flexible execution in arbitrary FL cross-silo settings. We built it on the open-source FL framework Flower [23]. Furthermore, we make this available as a training recipe with a fully transparent experimental configuration accompanying our results. We emphasize that federated generative pre-training can be done with affordable hardware configurations by interlinking single nodes containing $1$-$8$ GPUs from standard cloud providers —rather than renting multiple entire data centers concurrently. We also argue that federated systems can and must evolve beyond networks of clients holding both compute and data in a world of increasing hardware costs and competing privacy interests. To achieve this, Photon allows data providers (sources) to come together with compute providers in private partnerships amongst trusted parties without leaking any such data to other participants. We demonstrate that pre-training LLMs can be democratized through FL technologies, including data-rich actors independent of their connectivity and computing resources.
@@
Developing this novel federated system led to numerous insights, including the fact that, in defiance of expectations, larger federated LLMs can more easily find a consensus across clients than their smaller counterparts. Our results show that:
@@
    Under our system, Photon, larger models require less frequent communication while receiving a more significant boost in generalization performance than smaller ones.
@@
    The classical challenges of statistical heterogeneity and partial participation are much less impactful for overparameterized LLMs than the small models typically used in the FL literature. This opens the door to compute-efficient collaboration at a massive scale.
  4. 4.
    By utilizing our local execution engine, Pollen [24], to balance loads across computational nodes, we can establish a training pipeline that is faster and more robust than its centralized counterpart.
@@ 2 The Landscape of LLM Training
Generative pre-trained large language models (LLMs) have demonstrated powerful performance across various natural language processing tasks, leading to rapid and widespread adoption. They are trained on massive corpora of text, incorporating mixtures of low-quality web-scraped data and high-quality curated datasets [25, 26]. Recently, several institutions have released their pre-trained LLMs, either in the closed form, such as GPT-4 [26], Chinchilla [6], and Gemini [27], or the open-sourced form, such as BLOOM [28], LLaMa [29, 30], and Falcon [31]. The scaling laws identified by Kaplan et al. [2], Hoffmann et al. [6] dictate that model size and dataset size should be increased in equal measure to improve model performance best. These suggest a future race between entities interested in developing state-of-the-art LLMs to grab as many compute and data sources as possible. Thus, LLMs are headed in a promising direction that can become even more luminous by gaining the trust of private entities possessing an unprecedented breadth of knowledge and computing resources [16].
In Section 2.1, we describe the current landscape for generative pre-training of LLMs with particular attention to the techniques for centralized distributed training. We discuss federated learning (FL) and its impact as a communication efficient technique in Section 2.2. Section 2.3, for completeness, presents the effort the community has put into matching FL with LLMs fine-tuning. However, we highlight that our work tackles the far more challenging problem of federated pre-training of LLMs.
@@ 2.1 Centralized Distributed Optimization
LLM pre-training is based on two denoising pillars: leveraging huge batch sizes and very long contexts, i.e., sequence length of a single input sample. The combination of model and batch sizes forces training LLMs to scale SGD beyond the confines of a single GPU. Thus, it is often necessary (a) to process more training samples in parallel and (b) to split the model across GPUs if it cannot fit within the memory of one GPU. In the following, we present a subset of optimizations and techniques the research community proposed to make this challenging training recipe possible. As we will see in Section 2.2, FL follows as a natural next step in the sequence of distributed training optimizations previously adopted. Given the importance of these optimizations even for FL’s local training step, our work supports most of them, as described in Sections 5.1 and 6.1.
@@ 2.1.1 Data and Model Parallelism
The number of trainable parameters and the size of the datasets make LLM training very sensitive to the stochastic fluctuations of the optimizer used, thus requiring a solid and robust regularization achieved by the denoising properties of enormous batch sizes. Distributed Data Parallelism (DDP) replicates the model $N_{d}$ (number of devices) times across different devices to enable training with sufficiently large batch sizes. Then, DDP splits each enormous batch into micro-batches, subsequently fed into each replica in parallel. The number of samples in every micro-batch depends on the available memory in the hardware accelerator; as such, it is often referred to as device batch size. Thus, each replica produces the gradients of its training micro-batch locally and accumulates them via simple summation. This accumulation is eventually followed by a synchronization step where these gradients are averaged and applied to the model. Since the batch size of an ML algorithm plays a crucial role in its convergence [32, 33], this synchronization step happens after each replica has accumulated sufficient micro-batches to match the desired true batch size. Modern DDP implementations such as the one used by PyTorch Distributed [34] use the Ring AllReduce algorithm popularized by Horovod [35] to reduce the gradients across replicas. The algorithm is implemented with low-level collective communication libraries like Nvidia’s NCCL or Facebook’s Gloo. Ring AllReduce scales proportionally to the amount of data being transmitted and the latency of the slowest connection between workers in the ring, making it highly sensitive to the network topology connecting GPUs.
In the case of LLM training, the model size often mandates computing gradients over batch sizes larger than what GPUs can support, even with DDP. Thus, practitioners are forced to use a micro-batch size dictated by the number of training samples that fit in the GPU, which is, in turn, limited by the VRAM occupancy due to model parameters and activations. These factors result in a complex interplay between the model’s memory consumption and the training pipeline’s efficiency. The resulting memory-efficiency trade-off is even more complex when considering how many ways the models may be partitioned across GPUs to reduce their VRAM consumption.
For sufficiently large models, the parameters must be split across GPU workers so that they fit in VRAM. Traditionally, this was achieved via Model Parallelism (MP) [36, 37], splitting individual tensors and their computation vertically across GPUs and communicating as needed. While model parallelism can reduce the per-GPU memory requirements linearly by the model parallelism degree $N_{m}$ (number of model splits), it induces communication overheads that are tolerable in a single-machine context with high inter-GPU bandwidth but cannot scale to multi-machine contexts.
@@ 2.1.2 Fully Sharded Data Parallelism
An alternative approach is to shard the model into equally-sized units amongst GPUs, with units potentially containing multiple layers, and then materialize the units, as necessary, to compute the activations during the forward pass via collective communication. This form of fully-sharded data parallelism [38, 39] reduces memory consumption linearly in the data-parallelism degree $N_{d}$ while increasing communication by $1.5$$\times$ compared to standard DDP. It is also possible to combine this methodology with other techniques for reducing memory consumption, such as model parallelism, activation checkpointing [40], or CPU offloading [41]. Activation checkpointing functions by recomputing activations during the backward pass rather than saving them in memory. CPU offloading refers to offloading either data to system RAM or operations to the CPU.
Given the memory requirements, it is crucial to consider the minimum number of GPUs necessary to train one model with reasonable efficiency, e.g., without extreme CPU offloading. This estimate provides a lower bound on the hardware that an organization requires to participate in any distributed training of an LLM and is generally determined by the model size and the micro-batch size. The micro-batch size is also known as the device batch size since it is bounded by the number of samples that can fit on a GPU device together with the model. By accounting for each organization’s independent resources and manipulating the amount of local computation they do together with the micro-batch size, our method can relax the lower bound within reasonable limits, thus allowing even organizations with weak hardware to participate in distributed training.
@@ 2.1.3 Bottlenecks for generative pre-training of LLMs
Similarly, hardware accelerators with enough memory and throughput to support LLM training are scarce and increasingly unaffordable to anyone interested except for the best-funded organizations. Hundreds to thousands of such accelerators are required with extremely high monetary costs for training and inference [28, 43]. Moreover, such accelerators must be largely uniform in specifications to avoid stragglers and need to be extremely well-connected both within a computational node and across nodes due to the synchronization requirements of data parallel or fully shared data parallel SGD [34, 38, 39]. These constraints pose significant system challenges that can be mitigated by extremely expensive data center configurations or by developing heterogeneity-aware pipelines that exploit the resources available despite their heterogeneity. The difficulties described above scale with model size, as splitting the model across the memory of several GPUs further increases communication demands [38, 39] and the optimization gaps.
@@ 2.1.4 Mitigation of LLMs demands
The wider AI community has gone to great lengths to circumvent the high requirements of LLM utilization and thus expand the data pool we can draw upon. Such endeavors have widely focused on locally running efficient inference with pre-trained model weights [29, 30], quantization [44], or parameter-efficient fine-tuning [45]. The recent work proposes Petal [46], which enables wide-scale collaboration for inference and parameter-efficient fine-tuning over the Internet by joining the resources of multiple parties. Their work assumes that clients provide inference jobs or data for fine-tuning, and servers execute the LLM inference/fine-tuning in a distributed pipeline parallel fashion. Their system makes clients responsible for holding the trainable parameters to enable efficient fine-tuning, with servers merely running forward passes and returning gradients for the pre-trained weights they store.
We argue that while methods exploiting pre-trained weights are highly beneficial to the broader community, they do not resolve the bottleneck of pre-training. Thus, the community is bound to the decisions made by organizations capable of training LLMs regarding their structure and the data they use, and they may suffer downstream consequences for such reliance. Performance degradation may happen since the pre-trained model, further fine-tuned for a specific downstream task, is the dominant upper bound for the downstream model’s performances. Thus, we propose developing systems capable of the distributed pre-training of LLMs that can accommodate the variety of hardware and data available in the AI community. Nevertheless, this new paradigm may include training from scratch if the model’s size permits, starting from pre-initialized weights and retraining the entire model instead of only a limited subset of parameters.
@@ 2.2 Federated Learning and Local SGD
Traditional machine learning involves using a central server that hosts the machine learning models and all the data in one place. In contrast, when using federated learning (FL) [19] algorithms, client devices collaboratively learn a shared global model using their local compute and data.
FL aims to collaboratively learn a global model while keeping private data on the device. Such a training pipeline occurs over multiple communication rounds. During each round, a fraction of the clients are selected and receive the global model from the server. Those selected clients then perform local training with their local data before sending the updated models back to the central server. Finally, the central server aggregates these updates via averaging [19] into a pseudo-gradient. It then uses a federated optimizer [47] to update its model based on the pseudo-gradient, creating a new global model. Then, this three-stage process is repeated.
@@
Despite these advantages, FL comes with two major challenges in the form of data and systems heterogeneity [50]. Data heterogeneity refers to the tendency of naturally generated and partitioned data to fail the IID assumption, which is common in centralized ML optimization. Systems heterogeneity refers to the ability of client hardware to vary in terms of computational ability, communication efficiency, or availability [51]. Both forms of heterogeneity are highly relevant for the distributed training of LLMs. Data heterogeneity may arise from participants holding texts in different languages, belonging to various genres, or varying in complexity. Systems heterogeneity is primarily present in the computational ability of the GPUs held by a specific client, their VRAM, and number, as well as the communication efficiency of said client.